# De Finetti’s problem with fixed transaction costs and regime switching

### v1

@@ 3. Explicit characterization of $\mathcal{M}_{\zeta}$
where the function $\phi^{-1}_{0}$ is the well-defined inverse function of $\phi_{0}$ that is defined by
@@
Suppose that $\mu_{+}\leq 0$ and $\mu_{-}>0$. For the following mutually exclusive and collectively exhaustive cases, the set $\mathcal{M}_{\zeta}$ is singleton and can be characterized explicitly as follows.
@@
For later use, we introduce the following functions. Under Case (ii), $[x_{0},a_{7}]\ni x\mapsto(\hat{g}^{\prime})_{-}^{-1}(x):=\inf\{y\in[0,a];g^{\prime}(y)\geq g^{\prime}(x)\}$ and define the inverse function of $\left.g^{\prime}\right|_{[a,x_{0}]}$ as $(\hat{g}^{\prime})_{+}^{-1}:[g^{\prime}(x_{0}),g^{\prime}(a_{7})]\rightarrow[a,x_{0}]$ with $a_{7}:=\sup\{x>x_{0}:g^{\prime}(x)\leq g^{\prime}(a)\}$. Recall that $\phi_{0}$ is given by (3.8). Furthermore, put
@@
$\omega_{4}(x)$ :=$:=\begin{cases}\int_{(\hat{g}^{\prime})^{-1}_{+}(g^{\prime}(x))}^{x}\left(1-\frac{g^{\prime}(s)}{g^{\prime}(x)}\right)\operatorname{d}\!s,&x\in[x_{0},x_{3})\\ \int_{0}^{x}\left(1-\frac{g^{\prime}(s)}{g^{\prime}(x)}\right)\operatorname{d}\!s,&x\in[x_{3},\infty).\end{cases}$ (3.12)
@@
    $\mathcal{M}_{\zeta}=\begin{cases}\{(\hat{w}_{1},\hat{w}_{2})\},&\text{if }\beta\in B_{1}\cup B_{3},\\ \{(\tilde{w}_{1},\tilde{w}_{2})\},&\text{if }\beta\in B_{2}\cap\overline{B}_{3},\\ \{(\hat{w}_{1},\hat{w}_{2})\}\cup\{(\tilde{w}_{1},\tilde{w}_{2})\},&\text{otherwise,}\end{cases}$
    where, $(\tilde{w}_{1},\tilde{w}_{2}):=((\hat{g}^{\prime})^{-1}_{+}(g^{\prime}(\omega^{-1}_{4}(\beta))),\omega^{-1}_{4}(\beta))$, $(\hat{w}_{1},\hat{w}_{2}):=(0,\omega_{3}^{-1}(\beta))$, $B_{1}:=\{\beta>0:g^{\prime}(\omega^{-1}_{3}(\beta))<g^{\prime}(\omega^{-1}_{4}(\beta))\}$, $B_{2}:=\{\beta>0:g^{\prime}(\omega^{-1}_{3}(\beta))>g^{\prime}(\omega^{-1}_{4}(\beta))\}$, $B_{3}:=(\omega_{4}(x_{3}),\infty)$,  with $\omega^{-1}_{3}$ and $\omega^{-1}_{4}$ being the inverse functions of $\omega_{3}$ and $\omega_{4}$, respectively. Note that, when $\beta=\omega_{3}(x_{4})$, one has $\omega^{-1}_{3}(\beta)=\{(\hat{g}^{\prime})^{-1}_{-}(g^{\prime}(x_{4})),x_{4}\}$, in which case, $\{(\hat{w}_{1},\hat{w}_{2})\}$ is understood as $\{(0,(\hat{g}^{\prime})^{-1}_{-}(g^{\prime}(x_{4})),(0,x_{4})\}$.
@@ 4. Characterization of the optimal impulsive strategy
Let $(z_{1},z_{2})$ be an element of $\mathcal{M}_{\zeta}$. Then, the $(z_{1},z_{2})$-strategy is an optimal impulsive strategy to the control problem (2.4), if one of the following conditions holds true:
@@
Suppose $z_{2}\leq a.$ Using Propositions 3.1-3.4, and Theorems 3.1-3.4, one sees that $g^{\prime\prime}(x)>0$ for all $x\in[z_{2},a)$. Thus
@@
By (2.15), $V^{z_{2}}_{z_{1}}(x)\in C^{1}(\mathbb{R}_{+})\cap C^{2}(\mathbb{R}_{+}\backslash\{a,z_{2}\})$. We next verify $(\mathcal{A}-q)V^{z_{2}}_{z_{1}}(x)\leq 0$ for $x\in[0,\infty)\backslash\{a,z_{2}\}$. Using (2.11), (2.15), and the fact that $(\mathcal{A}-q)g^{\pm}(x)=0$ for all $x\neq a$, we have
@@
Together with (4.5), we proved $(\mathcal{A}-q)V^{z_{2}}_{z_{1}}(x)\leq 0$ on $(0,\infty)\backslash\{a,z_{2}\}$ when one of Conditions (a), (b), or (c) holds true. Combining with (4.1) and Lemma 2.2, we prove Theorem 4.1. ∎
@@
  • In Cases (i) and (ii), $\mathcal{M}_{\zeta}$ is a singleton set, and we have $z_{2}>a$.
  • In Case (iii), $\mathcal{M}_{\zeta}$ is a singleton set, and we have $g^{\prime\prime}(a+)>0$.
@@
    holds true, then $\mathcal{M}_{\zeta}$ is not necessarily a singleton set, but there is at least one $(z_{1},z_{2})\in\mathcal{M}_{\zeta}$ with $z_{2}>a$.
(2) Assume $\mu_{\pm}\leq 0$. Then $\mathcal{M}_{\zeta}$ is a singleton set, and we have $g^{\prime\prime}(a+)>0$.
(3) Assume $\mu_{+}\leq 0$ and $\mu_{-}>0$. Then $\mathcal{M}_{\zeta}$ is a singleton set.
@@
  • In Case (i), $\mathcal{M}_{\zeta}$ is a singleton set, and we have $g^{\prime\prime}(a+)>0$.
@@
[…]
    holds true, then $\mathcal{M}_{\zeta}$ may not be a singleton set, but there is at least one $(z_{1},z_{2})\in\mathcal{M}_{\zeta}$ with $z_{2}>a$.
@@ Appendix A Proofs
Combining (A.1)-(A.2) and using the arbitrariness of $\pi$ yields the upper bound.
@@
By (2.11) and the fact that $\𝑎g\in C^{1}(\mathbb{R})\cap C^{2}(\mathbb{R}\backslash\{a\})$, for any $x\in\mathbb{R}$, it holds that
@@
By (2.11) and (A.3), one knows that there exists $z_{1}^{0}\in(0,\infty)$ such that $g^{\prime}(z)$ is increasing over $[z_{1}^{0},\infty)$, which implies $\frac{\partial}{\partial z_{1}}\zeta(z_{1},z_{2})<0$ for all $z_{1}\geq z_{1}^{0}$. In addition, it holds that
@@
    $=(1-c_{-})e^{(\theta_{1}^{-}-\theta_{2}^{-})a}\left[(\theta_{2}^{-})^{2}e^{\theta_{2}^{-}x}-(\theta_{1}^{-})^{2}e^{-\theta_{1}^{-}x}\right],$
@@
    $h_{1}(a):=g^{+}(0)-c_{+}g^{-}(0)=e^{-\theta_{2}^{-}a}(1-c_{+}c_{-})-c_{+}(1-c_{-})e^{\theta_{1}^{-}a},$
    where we used (2.10) in the second equality. Due to the fact that $g^{-}(0)>g^{+}(0)>0$ (see (A.3)), if $c_{+}\leq 0$, we have $h_{1}(a)>0$ for all $a>0$. If $c_{+}>0$, it follows from $01-c_{+}c_{-}>$ 0 that the function $\mathbb{R}_{+}\ni x\mapsto h_{1}(x)$ is strictly decreasing and has a unique zero $a_{2}>0$ given by (3.2), where we have used the fact that $c_{+}>0$ implies $\theta_{2}^{+}-\theta_{2}^{-}>0$ (see (2.9)). Therefore, if $c_{+}>0$ and $a\in(0,a_{2})$, we have $h_{1}(a)>0$; and, if $c_{+}>0$ and $a\geq a_{2}$, we have $h_{1}(a)\leq 0$. In sum, we have
    $h_{1}(a)\begin{cases}>0\,\,\text{ if }c_{+}\leq 0,\\ >0\,\,\text{ if }c_{+}>0\text{ and }0<a<a_{2},\\ \leq 0\,\,\text{ if }c_{+}>0\text{ and }a\geq a_{2}.\end{cases}$
@@
        $-(1-c_{-})\Theta=-(1-c_{-})\left[c_{+}(\theta_{1}^{+})^{2}+(1-c_{+})(\theta_{2}^{+})^{2}\right]<0,$
@@
        Hence, the function $\mathbb{R}_{+}\ni x\mapsto h_{2}(x)$ is strictly decreasing and admits a unique zero $a_{3}>0$ given by (3.2). Hence, if $c_{+}>0$ and $0<a<a_{2}\leq a_{3}$, we have $-g^{\prime\prime}(a+)=h_{2}(a)>0$ on $a\in(0,a_{2})$, which implies $x_{0}>a$, and hence $g^{\prime\prime}(x)<0$ on $(a,x_{0})$ and $g^{\prime\prime}(x)>0$ on $(x_{0},\infty)$. Similarly, if $c_{+}>0$ and $0<a<a_{3}<a_{2}$, one knows that $-g^{\prime\prime}(a+)=h_{2}(a)>0$ and $x_{0}>a$, and hence $g^{\prime\prime}(x)<0$ on $(a,x_{0})$ and $g^{\prime\prime}(x)>0$ on $(x_{0},\infty)$.
@@
        If $c_{+}>0$ and $0<a_{3}\leq a\leq a_{2}$, then $g^{\prime\prime}(x)>0$ for all $x>a$. Actually, in the case $c_{+}>0$ and $0<a_{3}\leq a\leq a_{2}$, we get $-g^{\prime\prime}(a+)=h_{2}(a)\leq 0$, which means $x_{0}\leq a$.
@@
        If $\Theta\leq 0$, $c_{+}\leq 0$ and $a>0$, then $x_{0}>a$, $g^{\prime\prime}(x)<0$ on $(a,x_{0})$, and, $g^{\prime\prime}(x)>0$ on $(x_{0},\infty)$. Indeed, if $\Theta\leq 0$ and $a>0$, it follows from $h_{2}(0)>0$ that $-g^{\prime\prime}(a+)=h_{2}(a)>0$, which means $x_{0}>a$.
@@
        If $\Theta>0$, $c_{+}\leq 0$ and $a\in(0,a_{3})$, then $x_{0}>a$, $g^{\prime\prime}(x)<0$ on $(a,x_{0})$, and, $g^{\prime\prime}(x)>0$ on $(x_{0},\infty)$. Indeed, if $\Theta>0$ and $a\in(0,a_{3})$, it follows from $h_{2}(0)>0$ that the function $\mathbb{R}_{+}\ni x\mapsto h_{2}(x)$ is strictly decreasing and $a_{3}>0$ given by (3.2) is its unique zero. Hence, one knows that $-g^{\prime\prime}(a+)=h_{2}(a)>0$, which yields $x_{0}>a$.
@@
        If $\Theta>0$, $c_{+}\leq 0$ and $a\in[a_{3},\infty)$, then $g^{\prime\prime}(x)\geq 0$ for all $x>a$. Indeed, if $\Theta>0$ and $a\in[a_{3},\infty)$, one knows that $-g^{\prime\prime}(a+)=h_{2}(a)\leq 0$ for $a\in[a_{3},\infty)$, which means $x_{0}\leq a$.
@@
To start, we characterize the set $\mathcal{N}$ (Actually, if $\mathcal{N}$ is identified to be a singleton set, then by Proposition 2.2 and the relation $\mathcal{M}_{\zeta}\subseteq\mathcal{N}$ one has $\mathcal{M}_{\zeta}=\mathcal{N}$). For any $(z_{1},z_{2})\in\mathcal{N}$, either $z_{1}=0$ or $z_{1}>0$ holds true.
@@
    If there is a $(z_{1},z_{2})\in\mathcal{N}$ with $z_{1}>0$, then we have (2.18) and $g^{\prime}(z_{1})=g^{\prime}(z_{2})$, which forces us to conclude that $0<z_{1}\leq a\leq z_{2}<\inf\{x\geq a;g^{\prime}(x)\geq g^{\prime}(0)\}$. Then, it holds that $z_{1}=(g^{\prime})_{-}^{-1}(g^{\prime}(z_{1}))=(g^{\prime})_{-}^{-1}(g^{\prime}(z_{2}))$. Hence, (2.18) can be rewritten as
@@
    which inherits from $g^{\prime\prime}(x)$ the property of being positive on $(a,a_{4})$. That is to say, the unary function $\phi(x)$ defined by (3.3) is continuous and strictly increasing on $[a,a_{4}]$ with $\phi(a)=\psi((g^{\prime})_{-}^{-1}(g^{\prime}(a)),a)=\psi(a,a)=0$. Hence
@@
    If there is a $(z_{1},z_{2})\in\mathcal{N}$ with $z_{1}=0$, then, (2.14) holds true with $z_{1}=0$, that is
@@
Summing up the above results, we arrive at the following conclusion.
@@
    Due to the fact that
@@
    $\phi_{0}^{-1}(\beta)>\phi^{-1}(\beta).$
    Since both points of $\mathcal{N}$ are solutions to (2.14), by above and the fact that $g^{\prime}(x)$ is strictly increasing on $(a,a_{4})$, one can get
@@
$z^{\prime}_{2}<(g^{\prime})^{-1}_{2}(g^{\prime}(z_{2})).$

### v2

@@ 3. Explicit characterization of $\mathcal{M}_{\zeta}$
where the function $\phi^{-1}_{0}$ is the well-defined inverse function of $\phi_{0}$ given by
@@
Suppose that $\mu_{+}\leq 0$ and $\mu_{-}>0$. For the following mutually exclusive and collectively exhaustive cases, the set $\mathcal{M}_{\zeta}$ is a singleton set and can be characterized explicitly as follows.
@@
For later use, we introduce the following functions. Under Case (ii), $x_{0},a_{7}]\ni x\mapsto(\hat{g}^{\prime})_{-}^{-1}(x):=\inf\{y\in[0,a];g^{\prime}(y)\geq g^{\prime}(x)\}$ and define the inverse function of $\left.g^{\prime}\right|_{[a,x_{0}]}$ as $(\hat{g}^{\prime})_{+}^{-1}:[g^{\prime}(x_{0}),g^{\prime}(a_{7})]\rightarrow[a,x_{0}]$ with $a_{7}:=\sup\{x>x_{0}:g^{\prime}(x)\leq g^{\prime}(a)\}$. Recall that $\phi_{0}$ is given by (3.8). Furthermore, put
@@
$\omega_{4}(x)$ $:=\begin{cases}\int_{(\hat{g}^{\prime})^{-1}_{+}(g^{\prime}(x))}^{x}\left(1-\frac{g^{\prime}(s)}{g^{\prime}(x)}\right)\operatorname{d}\!s,&x\in[x_{0},x_{3})\\ \int_{0}^{x}\left(1-\frac{g^{\prime}(s)}{g^{\prime}(x)}\right)\operatorname{d}\!s,&x\in[x_{3},\infty).\end{cases}$ (3.12)
@@
    $\mathcal{M}_{\zeta}=\begin{cases}\{(\hat{w}_{1},\hat{w}_{2})\},&\text{if }\beta\in B_{1}\cup B_{3},\\ \{(\tilde{w}_{1},\tilde{w}_{2})\},&\text{if }\beta\in B_{2}\cap\overline{B}_{3},\\ \{(\hat{w}_{1},\hat{w}_{2})\}\cup\{(\tilde{w}_{1},\tilde{w}_{2})\},&\text{if }\beta\in\overline{B_{1}\cup B_{2}\cup B_{3}},\end{cases}$
    where, $(\tilde{w}_{1},\tilde{w}_{2}):=((\hat{g}^{\prime})^{-1}_{+}(g^{\prime}(\omega^{-1}_{4}(\beta))),\omega^{-1}_{4}(\beta))$, $(\hat{w}_{1},\hat{w}_{2}):=(0,\omega_{3}^{-1}(\beta))$, $B_{1}:=\{\beta>0:g^{\prime}(\omega^{-1}_{3}(\beta))<g^{\prime}(\omega^{-1}_{4}(\beta))\}$, $B_{2}:=\{\beta>0:g^{\prime}(\omega^{-1}_{3}(\beta))>g^{\prime}(\omega^{-1}_{4}(\beta))\}$, $B_{3}:=(\omega_{4}(x_{3}),\infty)$,  with $\omega^{-1}_{3}$ and $\omega^{-1}_{4}$ being the inverse functions of $\omega_{3}$ and $\omega_{4}$, respectively. It holds that $B_{2}\cap B_{3}=\emptyset$. We mention that, when $\beta=\omega_{3}(x_{4})$, one has $\omega^{-1}_{3}(\beta)=\{(\hat{g}^{\prime})^{-1}_{-}(g^{\prime}(x_{4})),x_{4}\}$; in this case, $\{(\hat{w}_{1},\hat{w}_{2})\}$ is understood as $\{(0,(\hat{g}^{\prime})^{-1}_{-}(g^{\prime}(\omega^{-1}_{3}(\beta))),(0,x_{4})\}$.
@@ 4. Characterization of the optimal impulsive strategy
Let $(z_{1},z_{2})$ be the element of $\mathcal{M}_{\zeta}$. Then, the $(z_{1},z_{2})$-strategy is an optimal impulsive strategy to the control problem (2.4), if one of the following conditions holds true:
@@
Suppose $z_{2}\leq a.$ Using Propositions 3.1-3.4, and Theorems 3.1-3.4, one sees that $g^{\prime\prime}(x)>0$ for all $x\in[z_{2},a)$.
@@
[…]
@@
[…]
@@
  • In Cases (i) and (ii), $\mathcal{M}_{\zeta}$ is a singleton, and we have $z_{2}>a$.
  • In Case (iii), $\mathcal{M}_{\zeta}$ is a singleton, and we have $g^{\prime\prime}(a+)>0$.
@@
    holds true, then $\mathcal{M}_{\zeta}$ is not necessarily a singleton, but there is at least one $(z_{1},z_{2})\in\mathcal{M}_{\zeta}$ with $z_{2}>a$.
(2) Assume $\mu_{\pm}\leq 0$. Then $\mathcal{M}_{\zeta}$ is a singleton, and we have $g^{\prime\prime}(a+)>0$.
(3) Assume $\mu_{+}\leq 0$ and $\mu_{-}>0$. Then $\mathcal{M}_{\zeta}$ is a singleton.
@@
  • In Case (i), $\mathcal{M}_{\zeta}$ is a singleton, and we have $g^{\prime\prime}(a+)>0$.
@@
    • $\beta<\omega_{3}(x_{4})$ and $g^{\prime}(\omega^{-1}_{3}(\beta))\geq g^{\prime}(\omega^{-1}_{4}(\beta))$,
    holds true, then $\mathcal{M}_{\zeta}$ may not be a singleton, but there is at least one $(z_{1},z_{2})\in\mathcal{M}_{\zeta}$ with $z_{2}>a$.
@@
5. Numerical Analysis and Economic Interpretations
To complement the theoretical results derived in the previous sections, we now present a series of numerical experiments. These analyses serve to visualize the optimal dividend strategy in action and to provide economic intuition for how the firm’s decisions are shaped by the underlying model parameters.
5.1. A sample path under the optimal strategy
We begin by simulating a sample path of the optimal controlled surplus process $U^{\pi^{*}}$. The left panel of Figure 1 illustrates this path over a time horizon of $T=10$. The baseline parameters for this simulation are set as follows: the regime-switching threshold is $a=1$; the drift and volatility coefficients are $(\mu_{-},\sigma_{-})=(0.5,0.5)$ for the lower regime $(U_{t}\leq a)$ and $(\mu_{+},\sigma_{+})=(0.1,0.1)$ for the upper regime $(U_{t}>a)$; and the fixed transaction cost $\beta=0.5$. For this parameter set, the explicit characterization of $\mathcal{M}_{\zeta}$ derived in Section 3 identifies the optimal dividend barriers as $z^{*}_{1}=0.4277$ and $z^{*}_{2}=1.9059$. This strategy constitutes the optimal dividend strategy, since Condition $(a)$ of Theorem 4.1 holds. This models a company with a high-growth, high-risk “startup” phase (when surplus is below $a$) that transitions into a low-growth, low-risk “maturity” phase upon expansion (surplus above $a$). As observed in the left panel of Figure 1, the firm sets a relatively low upper barrier $z^{*}_{2}$. It allows the surplus to enter the more stable (but less profitable) mature phase to safely accumulate funds and pay a dividend of size $\Delta L_{t}=z^{*}_{2}-z^{*}_{1}$. The dividend resets the surplus to $z^{*}_{1}$, positioning the firm to re-leverage its high-growth startup phase. The strategy is thus a dynamic cycle of navigating between risk and stability.
The right panel of Figure 1 represents a completely different economic reality over a time horizon of $T=50$. Here, the parameters are reversed: $(\mu_{+},\sigma_{+})=(0.5,0.5)$ and $(\mu_{-},\sigma_{-})=(0.1,0.1)$ with $\beta=0.5$ and $a=1$. This models a company that is stable but stagnant when small, and only enters a high-growth, high-risk expansion phase after its surplus exceeds the threshold $a$. The firm endures a long period of slow growth below $a$, as the low drift makes it difficult to cross the threshold. Once the threshold is crossed, the firm enters the highly profitable expansion phase. The optimal strategy is to establish a high upper barrier at $z_{2}^{*}=10.4512$, a value consistent with Condition $(a)$ of Theorem 4.1, to capitalize on this phase. Upon reaching $z^{*}_{2}$, the firm pays out a very large dividend of $\Delta L_{t}=z^{*}_{2}-z^{*}_{1}$. The lower barrier is set at $z^{*}_{1}=0$, the ruin level itself. It signifies that the optimal path is not to restart, but to perform a terminal payout. This aggressive strategy is rational for shareholders aiming to fully extract the firm’s value after a successful high-growth period, rather than risk returning to the stagnant, low-growth phase.
Figure 1. A sample path under the optimal strategy
5.2. Sensitivity analysis of model parameters
We now investigate the sensitivity of the optimal dividend barriers $(z^{*}_{1},z^{*}_{2})$ with respect to the key model parameters: the transaction cost $\beta$, the regime-switching threshold $a$, the lower regime drift $\mu_{-}$, and the lower regime volatility $\sigma_{-}$.
The top-left panel of Figure 2 shows the effect of the transaction cost $\beta$ on the optimal barriers under $(\mu_{+},\sigma_{+})=(0.1,0.1)$, $(\mu_{-},\sigma_{-})=(0.5,0.5)$ and $a=1$. The results are in complete agreement with Propositions 3.5 and 3.6, as well as satisfying Condition $(a)$ of Theorem 4.1. As the cost $\beta$ of paying a dividend increases, the upper barrier $z^{*}_{2}$ increases, while the lower barrier $z^{*}_{1}$ decreases. Consequently, the dividend size $\Delta L_{t}=z^{*}_{2}-z^{*}_{1}$ becomes a strictly increasing function of the cost. The intuition is straightforward: a higher fixed cost per transaction incentivizes the firm to pay dividends less frequently but in larger amounts.
We now examine the sensitivity of the optimal strategy to the regime-switching threshold $a$ under $(\mu_{+},\sigma_{+})=(0.5,0.1)$, $(\mu_{-},\sigma_{-})=(1,0.5)$ and $\beta=1$. Each $z^{*}_{2}$ value shown in the figure satisfies Condition $(a)$ of Theorem 4.1. When $a$ is small, the high-growth zone is narrow, forcing the firm to set a high $z^{*}_{2}$ to patiently accumulate surplus in the safe upper regime. As $a$ increases, making the high-growth zone more accessible, the strategy becomes more aggressive by progressively lowering $z^{*}_{2}$ to realize profits sooner. The sharp kink at $a=2.6$ marks a fundamental strategic shift. At and beyond this point, the firm prefers to pay a dividend precisely when it is about to be pushed into the low-growth zone, immediately resetting the process to stay within its preferred, more profitable environment.
The bottom-left panel illustrates the impact of the lower regime’s profitability $\mu_{-}$ under $(\mu_{+},\sigma_{+})=(0.5,1)$, $\sigma_{-}=0.5$, $a=2$ and $\beta=1$. Each $z^{*}_{2}$ value shown in the figure satisfies Condition (a) of Theorem 4.1. When $\mu_{-}$ is negative, the lower regime is a “danger zone.” The firm adopts a highly conservative strategy, maintaining high barriers $z_{1}^{*}=4,z_{2}^{*}=10$ to create a large safety buffer and minimize the risk of ruin after a dividend payout. As $\mu_{-}$ increases and surpasses the upper regime’s drift $\mu_{+}$, the lower regime becomes the engine of growth. The firm’s strategy flips to become aggressive. It sets both barriers significantly lower to ensure that after a payout, the process returns to the highly profitable lower regime, and it pays dividends more quickly to minimize time spent in the less profitable upper regime.
Finally, the bottom-right panel shows the effect of the lower regime’s risk $\sigma_{-}$ under $(\mu_{+},\sigma_{+})=(0.5,0.5)$, $\mu_{-}=1$, $\beta=1$ and $a=8$. Both barriers $z^{*}_{1}$ and $z^{*}_{2}$ are monotonically increasing functions of $\sigma_{-}$. For $\sigma_{-}>0.63$, $z_{2}^{*}$ satisfies condition $(a)$ of Theorem 4.1; for $\sigma_{-}<0.63$, both Conditions $(b)$ and $(c)$ hold true. This reflects the firm’s response to an increase in operational risk. A higher volatility in the primary operating regime increases the probability of ruin. To mitigate this risk, the firm adopts a more conservative policy by holding more precautionary cash. It raises $z^{*}_{1}$ to leave a larger buffer and consequently must also raise $z^{*}_{2}$ to ensure that the payout remains large enough to justify the transaction cost. Higher risk, therefore, leads to delayed and larger dividend payments.
Figure 2. Sensitive of parameters $\beta$, $a$, $\mu_{-}$ and $\sigma_{-}$
@@ Appendix A Proofs
Combining (A.1)-(A.2) and using the arbitrariness of $\pi$ yields the upper bound of (2.5).
@@
By (2.13) and the fact that $g\in C^{1}(\mathbb{R})\cap C^{2}(\mathbb{R}\backslash\{a\})$, for any $x\in\mathbb{R}$, it holds that
@@
By (2.13) and (A.3), one knows that there exists $z_{1}^{0}\in(0,\infty)$ such that $g^{\prime}(z)$ is increasing over $z_{1}^{0},\infty)$, which implies $\frac{\partial}{\partial z_{1}}\zeta(z_{1},z_{2})<0$ for all $z_{1}\geq z_{1}^{0}$. In addition, it holds that
@@
    $=(1-c_{-})e^{(\theta_{1}^{-}-\theta_{2}^{-})a}\left[(\theta_{2}^{-})^{2}e^{\theta_{2}^{-}x}-(\theta_{1}^{-})^{2}e^{-\theta_{1}^{-}x}\right],$ (A.7)
@@
    $h_{1}(a):=g^{+}(0)-c_{+}g^{-}(0)=e^{-\theta_{2}^{-}a}(1-c_{+}c_{-})-c_{+}(1-c_{-})e^{\theta_{1}^{-}a},$ (A.9)
    where we used (2.12) in the second equality. Due to the fact that $g^{-}(0)>g^{+}(0)>0$ (see (A.3)), if $c_{+}\leq 0$, we have $h_{1}(a)>0$ for all $a>0$. If $c_{+}>0$, it follows from $1-c_{+}c_{-}>0$ that the function $\mathbb{R}_{+}\ni x\mapsto h_{1}(x)$ is strictly decreasing and has a unique zero $a_{2}>0$ given by (3.2), where we have used the fact that $c_{+}>0$ implies $\theta_{2}^{+}-\theta_{2}^{-}>0$ (see (2.11)). Therefore, if $c_{+}>0$ and $a\in(0,a_{2})$, we have $h_{1}(a)>0$; and, if $c_{+}>0$ and $a\geq a_{2}$, we have $h_{1}(a)\leq 0$. In sum, we have
[…]
@@
        $(1-c_{-})\Theta=-(1-c_{-})\left[c_{+}(\theta_{1}^{+})^{2}+(1-c_{+})(\theta_{2}^{+})^{2}\right]<0,$
@@
        Hence, the function $\mathbb{R}_{+}\ni x\mapsto h_{2}(x)$ is strictly decreasing and admits a unique zero $a_{3}>0$ given by (3.2). Hence, if $c_{+}>0$ and $0<a<a_{2}\leq a_{3}$, we have $g^{\prime\prime}(a+)=h_{2}(a)>0$ on $a\in(0,a_{2})$, which implies $x_{0}>a$, and hence $g^{\prime\prime}(x)<0$ on $(a,x_{0})$ and $g^{\prime\prime}(x)>0$ on $(x_{0},\infty)$. Similarly, if $c_{+}>0$ and $0<a<a_{3}<a_{2}$, one knows that $g^{\prime\prime}(a+)=h_{2}(a)>0$ and $x_{0}>a$, and hence $g^{\prime\prime}(x)<0$ on $(a,x_{0})$ and $g^{\prime\prime}(x)>0$ on $(x_{0},\infty)$.
@@
        If $c_{+}>0$ and $0<a_{3}\leq a\leq a_{2}$, then $g^{\prime\prime}(x)>0$ for all $x>a$. Actually, in the case $c_{+}>0$ and $0<a_{3}\leq a\leq a_{2}$, we get $g^{\prime\prime}(a+)=h_{2}(a)\leq 0$, which means $x_{0}\leq a$.
@@
        If $\Theta\leq 0$, $c_{+}\leq 0$ and $a>0$, then $x_{0}>a$, $g^{\prime\prime}(x)<0$ on $(a,x_{0})$, and, $g^{\prime\prime}(x)>0$ on $(x_{0},\infty)$. Indeed, if $\Theta\leq 0$ and $a>0$, it follows from $h_{2}(0)>0$ that $g^{\prime\prime}(a+)=h_{2}(a)>0$, which means $x_{0}>a$.
@@
        If $\Theta>0$, $c_{+}\leq 0$ and $a\in(0,a_{3})$, then $x_{0}>a$, $g^{\prime\prime}(x)<0$ on $(a,x_{0})$, and, $g^{\prime\prime}(x)>0$ on $(x_{0},\infty)$. Indeed, if $\Theta>0$ and $a\in(0,a_{3})$, it follows from $h_{2}(0)>0$ that the function $\mathbb{R}_{+}\ni x\mapsto h_{2}(x)$ is strictly decreasing and $a_{3}>0$ given by (3.2) is its unique zero. Hence, one knows that $g^{\prime\prime}(a+)=h_{2}(a)>0$, which yields $x_{0}>a$.
@@
        If $\Theta>0$, $c_{+}\leq 0$ and $a\in[a_{3},\infty)$, then $g^{\prime\prime}(x)\geq 0$ for all $x>a$. Indeed, if $\Theta>0$ and $a\in[a_{3},\infty)$, one knows that $g^{\prime\prime}(a+)=h_{2}(a)\leq 0$ for $a\in[a_{3},\infty)$, which means $x_{0}\leq a$.
@@
To start, we characterize the set $\mathcal{N}$ (Actually, if $\mathcal{N}$ is identified to be a singleton, then by Proposition 2.2 and the relation $\mathcal{M}_{\zeta}\subseteq\mathcal{N}$ one has $\mathcal{M}_{\zeta}=\mathcal{N}$). For any $(z_{1},z_{2})\in\mathcal{N}$, either $z_{1}=0$ or $z_{1}>0$ holds true.
@@
    We first check whether or not there is a $(z_{1},z_{2})\in\mathcal{N}$ such that $z_{1}>0$. If there is a $(z_{1},z_{2})\in\mathcal{N}$ with $z_{1}>0$, then we have (2.22) and $g^{\prime}(z_{1})=g^{\prime}(z_{2})$, which forces us to conclude that $0<z_{1}\leq a\leq z_{2}<\inf\{x\geq a;g^{\prime}(x)\geq g^{\prime}(0)\}$. Then, it holds that $z_{1}=(g^{\prime})_{-}^{-1}(g^{\prime}(z_{1}))=(g^{\prime})_{-}^{-1}(g^{\prime}(z_{2}))$. Hence, (2.22) can be rewritten as
@@
    which inherits from $g^{\prime\prime}(x)$ the property of being positive on $(a,a_{4})$. That is to say, the unary function $\phi(x)$ defined by (3.3) is continuous and strictly increasing on $a,a_{4}]$ with $\phi(a)=\psi((g^{\prime})_{-}^{-1}(g^{\prime}(a)),a)=\psi(a,a)=0$. Hence
@@
    We then check whether or not there is a $(z_{1},z_{2})\in\mathcal{N}$ such that $z_{1}=0$. If there is a $(z_{1},z_{2})\in\mathcal{N}$ with $z_{1}=0$, then, (2.18) holds true with $z_{1}=0$, that is
@@
Summing up the above arguments, we arrive at our conclusion as follows.
@@
    Here, $\phi_{0}^{-1}$ denotes the well-defined inverse function of $\phi_{0}$. Due to the fact that
@@
    $\phi_{0}^{-1}(\beta)>\phi^{-1}(\beta).$ (A.13)
    Since both points of $\mathcal{N}$ are solutions to (2.18), by (A.13) and the fact that $g^{\prime}(x)$ is strictly increasing on $(a,a_{4})$, one can get
@@
$z^{\prime}_{2}<(g^{\prime})^{-1}_{2}(g^{\prime}(z_{2})).$ (A.15)